# QueSTMaps: Queryable Semantic Topological Maps for 3D Scene Understanding

### v1

@@ IV-B Room Labeling
Baseline Methods. To benchmark our room labeling model separately, we compare it against four baseline methods, each representing different strategies within the domain. As the previous works used ground-truth object-to-room association, all baselines in this study are given the same privilege to ensure a fair comparison.
@@
    Zero-Shot LLM Baseline: It utilizes a generative language model [49] to generate descriptive strings about room compositions and applies a scoring method to identify the most appropriate room label [17].
@@
Results. As shown in Table Table II, our proposed CLIP Self-Attention Transformer model outperforms the next best algorithm, the embedding-based fine-tuning approach, by 10-12% in both F1 and mAP metrics. This improvement is significant, especially for room types with fewer examples, addressing the challenge of long-tail distribution. Using the self-attention transformer instead of averaging the object embeddings shows approximately a 12% improvement on the F1 and mAP metric. This specifically highlights the utility of our model in finding the room label-aligned embedding. Although the contrastive learning and logits prediction methods perform similarly, we continue to employ the contrastive learning framework for further experiments due to its compatibility with open-vocabulary room-type searches, facilitated by the room label-aligned CLIP embeddings.
@@ IV-C Complete Pipeline
We evaluate our end-to-end pipeline on the Matterport3D dataset. We show the results for room segmentation and labeling in Table Table III and compare it against the rich semantic results from RoomFormer [16], which is currently the state-of-the-art in room segmentation and labeling. Using our room and transition region outputs and room label outputs, we also construct semantic topological maps as shown in Fig. Figure 4.
To demonstrate the difference between the output of our method and other open-vocabulary 3D scene-understanding methods, we show a qualitative comparison in Fig. Figure 5. The comparison with OpenScene [8] shows that as the model computes per-point features, it is unable to correctly highlight the dense 3D region for a specific room type. We also contrast our results with the output from the 3D instance segmentation method OpenMask3D [12], to show that it cannot highlight dense regions given a room type. Although it can correctly identify objects for simpler queries like bedroom and kitchen, it struggles with concepts like living room and study room. These results clearly highlight the difference in the room-level understanding of 3D scenes between previously published methods and our proposed pipeline.
@@ V Limitations and Conclusion
Extensive experimention revealed that in cases where the room was not visible from a top-down view (due to noise or due to the room being very small), the room segmentation pipeline was unable to recognise it. In cases where there were no objects identified in a room by the instance mapping pipeline, our labeling model was unable to generate a CLIP feature for the room. Finally, some indoor environments such as open plan houses are better organised into “zones” rather than “rooms”. Since we rely on walls to segment our indoor scene, we cannot identify sub-room “zones”.
To summarise, our model extracts a natural-language queryable semantic topological map of 3D indoor scenes. We utilise a 3 channel Mask RCNN-based segmentation model to identify rooms and transition regions. We then utilise a custom self-attention transformer to generate a room-label aligned CLIP features for our rooms based on their confined objects. We outperform the state-of-the-art floorplan segmentation and labeling models on two challenging datasets. Additionally, we enable open-vocabulary room-level natural language queries on the scene. Future work may explore methods to segment a scene into zones. Language-queryable topological segmentation of a 3D scene into rooms can enable a robot to better reason about localization and enhance downstream tasks like planning [53] and navigation [39]; integration of our proposed method in such pipelines will be an exciting future work.

### v2

@@ IV-B Room Labeling
Baseline Methods. To benchmark our room labeling model separately, we compare it against different strategies within the domain. As the previous works used ground-truth object-to-room association, all baselines in this study are given the same privilege to ensure a fair comparison.
@@
    Zero-Shot LLM Baseline: Utilizes  [36] to generate descriptive strings about room compositions and applies a scoring method to identify the most appropriate room label [14].
@@
Results. As shown in Table Table II, our proposed CLIP Self-Attention Transformer model outperforms the next best algorithm, the embedding-based fine-tuning approach, by 10-12% in both F1 and mAP metrics. This improvement is significant, especially for room types with fewer examples, addressing the challenge of long-tail distribution. This specifically highlights the utility of our model in finding the room label-aligned embedding.
@@ IV-C Complete Pipeline
We evaluate our end-to-end pipeline on the Matterport3D dataset. We show the results for room segmentation and labeling in Table Table III and compare it against the rich semantic results from RoomFormer [13], the current SoTA in room segmentation and labeling. To demonstrate the difference between the output of our method and other open-vocabulary 3D scene-understanding methods, we show a qualitative comparison in Fig. Figure 4. The comparison with OpenScene [7] shows that as the model computes per-point features, it is unable to correctly highlight the dense 3D region for a specific room type. We also contrast our results with the output from the 3D instance segmentation method OpenMask3D [9], to show that it cannot highlight dense regions given a room type. Although it can correctly identify objects for simpler queries like bedroom and kitchen, it struggles with concepts like living room and study room. These results clearly highlight the difference in the room-level understanding of 3D scenes between previously published methods and our proposed pipeline. Moreover, our room and transition region outputs and room label outputs, enable the downstream applications such as natural-language laden instructions for navigation.
@@ V Limitations and Conclusion
In cases where the room was not visible from a top-down view (due to noise or due to the room being very small), the room segmentation pipeline was unable to recognise it. In cases where there were no objects identified in a room by the instance mapping pipeline, the labeling model was unable to generate a CLIP vector for the room. Some environments such as open plan houses are better organised into “zones” rather than “rooms”. Since we rely on walls to segment, we cannot identify sub-room “zones”.
To summarise, our model extracts a natural-language queryable semantic topological map of 3D indoor scenes. We utilise a 3 channel Mask RCNN-based segmentation model to identify rooms and transition regions. We then utilise a self-attention transformer to generate room-label aligned CLIP features for our rooms based on their contained objects. We outperform the SoTA floorplan segmentation and labeling models on two challenging datasets. Additionally, we enable open-vocabulary room-level natural language queries on the scene. Future work may explore methods to segment a scene into zones. Language-queryable topological segmentation of a 3D scene into rooms can enable a robot to better reason about localization and enhance downstream tasks like planning [40] and navigation [32]; integration of our method in such pipelines will be an exciting future work.